OpenStudy (anonymous):

4(p-10+8=-10 identify as a conditional a condratiction or a identity.

OpenStudy (anonymous):

4(p-1)+8=-10

OpenStudy (anonymous):

hey your back can u help me again

OpenStudy (anonymous):

What do you get when you simplify the equation?

OpenStudy (anonymous):

its the second one not the first one

OpenStudy (anonymous):

Yeah, I figured.

OpenStudy (anonymous):

So distribute the 4, and isolate your variable (if possible). What have ya got?

OpenStudy (anonymous):

use the disturbutive property...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

4*p 4*1?

OpenStudy (anonymous):

4(p-1) = 4p - 4 So.. \(4(p-1) + 8 =-10\) \[\implies 4p - 4 + 8 = -10\] Then we can combine the -4 and the +8

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

So what do you have now?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Please write the whole thing.

OpenStudy (anonymous):

4p-4=-8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That -8 should be a -10

OpenStudy (anonymous):

Otherwise, yes. That's correct. So now we can either keep simplifying it, or you can see if you can tell yet. Is this equation: Always true? Always false? Or does it depend on the variables?

OpenStudy (anonymous):

can we keep simplyfying it im still alittle lost

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

Sure, that's fine.

OpenStudy (anonymous):

Where did you feel like you got lost?

OpenStudy (anonymous):

why was it -10?

OpenStudy (anonymous):

Because we didn't change that side from the original equation.

OpenStudy (anonymous):

You had \[4(p-1) + 8 = -10\] After you distribute you have: \[4p-4 + 8 = -10\] Then you combine the +8 and the -4 and you get: \[4p - 4 = -10\] Then you can add 4 to both sides: 4p - 4 = -10 + 4 + 4 ------------ 4p = -6

OpenStudy (anonymous):

so its conditional

OpenStudy (anonymous):

It is! Nicely done.

OpenStudy (anonymous):

thanx to you11111!!!!!! now you have to teach me the solving part its really hard to me!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

thanx to you11111!!!!!! now you have to teach me the solving part its really hard to me!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Ok, I'm happy to help you practice if you like.

OpenStudy (anonymous):

If you like you can email me and we can arrange a time to practice solving/simplifying equations. Online whiteboards tend to work a bit better than this interface.